Data for Mercy Inc is as follows;

FCF= $10 million

WACC= 10%

\FCF is expected to grow at a constant rate of gL=5%

ST Investments = $20 million

Debt= $40 million

Preferred Stock= $10 million

Number of shares = 100,000

Mercy Inc, projected free cash flows are as follows

Year 1 FCF = $1 million

Year 2 FCF = $5 million

Year 3 FCF =$10 million

FCF grows at a constant rate of 5% after Year 3.

Question 1

(a) Assuming the WACC, ST Investment,Debt,Preferred Stock and number of shares from the data above,Calculate the expected Horizon Value of operations at Year 3?

(b) What is the Value of Operations for Mercy Inc based on the FCF projections above ?

(c) What is the Intrinsic Value of Equity to Common Stockholders and the Intrinsic Price per Share for Mercy Inc, based both sets of data above ?
Answer #1

a). Horizon Value = [FCF(1 + g)] / [WACC - g]

= [10 * (1 + 0.05)] / [0.10 - 0.05] = 10.5 / 0.05 = $210 million

b). Value of Operations = [FCF1 / (1 + WACC)] + [FCF2 / (1 + WACC)2] + [(FCF3 + Horizon Value) / {(WACC - g) * (1 + WACC)3]

= [1 / (1 + 0.10)] + [5 / (1 + 0.10)2] + [(10 + 210) / (1 + 0.10)3]

= 0.91 + 4.13 + 165.29 = $170.33 million

c). Intrinsic Value of Equity = Value of Operations - Debt - Preferred Stock + Short-term Investment

= 170.33 - 40 - 10 + 20 = $140.33 million

Intrinsic Price per Share = $140.33 million / Shares Outstanding

= 140.33 / 0.1 = $1,403.31